# A fast neural model for Bayesian Mixed-Effects Regression

### v1

@@ 4 Discussion
In this paper we present metabeta, a probabilistic transformer-based model that performs efficient approximate Bayesian inference for mixed-effects regression. We trained metabeta on simulated datasets with varying ranges for predictors, regression parameters, and outcomes. Most importantly, these datasets incorporate varying priors and we condition the model outputs on them, which not only amortizes the high computational costs encountered when using MCMC for parameter estimation, but also generalizes previous neural posterior estimation (NPE) techniques that are trained on a fixed prior. We show that our model has favorable and robust performance on in-distribution and out-of-distribution test sets, based on real hierarchical datasets practitioners care about. In each experiment, we compare the results of our model with Hamiltonian Monte Carlo (HMC), the gold-standard MCMC method for Bayesian mixed-effects regression, and show that metabeta often outperforms HMC in terms of accuracy, stability, and fidelity of credible intervals – all at a fraction of the time required for parameter estimation with HMC.
@@ 4.1 Limitations and Outlook
Each trained version of metabeta is currently tailored to the size of the regression problem in terms of the number of fixed effects ($d$) and the number of random effects ($q$). This means there are potentially as many model versions as combinations of $d$ and $q$. A single model with an upper bound to $d$ and $q$ is possible with our setup, but will never perform as well as a model that is specialized for a given problem size. Since a single snapshot of the weights requires about 2020mb of storage, pulling each model from an online server on-demand seems to be a manageable alternative to a single model with on average worse performance.
Our choice of model architecture trades of posterior expressivity for computation speed: Other normalizing flow methods like Neural Spline Flows (Durkan et al., 2019), Flow Matching (Wildberger et al., 2023), Conditional Diffusions (Chen et al., 2025; Reuter et al., 2025) or TarFlow (Zhai et al., 2025) offer more flexible posterior shapes, but posterior sampling is considerably more expensive than for Affine Coupling Flows. The relative simplicity of affine coupling posteriors can be seen as implicit regularization, preventing overly irregular quantification of regression parameter uncertainty.
A noteworthy outlook for metabeta is the incorporation of attention over predictors, in addition to the already present attention over samples (Müller et al., 2021). While this may introduce significant overhead for the summary network, it promises to handle dependencies between predictors in a more explicit and robust manner.
Finally, NPE models generally suffer when test data are strongly out of distribution in comparison to training data. While there are ways to amend this (Ward et al., 2022), a future iteration of metabeta will be trained directly on realistic hierarchical datasets generated by an LLM (Borisov et al., 2022; Wang et al., 2024; Jagadish et al., 2024; 2025).
@@ 4.2 Conclusion
Our model brings Bayesian mixed-effects regression closer to practical usability in real-world applications. In its current form, it already enables rapid prototyping – practitioners can quickly test different model specifications and validate findings using conventional tools if needed. Our analyses highlight that metabeta is immediately applicable to such use cases. Looking ahead, we envision scaling our model to larger regression problems, incorporating attention over predictors and training on more realistic, LLM-generated hierarchical data. This would open the door to entirely new applications of Bayesian mixed-effects regression that are currently out of reach.
@@ Appendix A Dataset Simulation
A.1 Predictors
We sample $n_{i}$ observations of predictor $j$, $\mathbf{x_{ij}}$, from the following distributions: Normal, StudentT, Uniform, Bernoulli, NegativeBinomial, ScaledBeta; with respective probability $(0.10,0.40,0.05,0.25,0.10,0.10)$, making student $t$ distributed predictors the most likely ones. For each dataset, the parameters of these distributions are randomly sampled with the constraint that the resulting outcome $\mathbf{y}$ has a bounded variance (i.e. does not explode). Correlation is induced by sampling $\mathbf{LL}^{\top}=\mathbf{R}\sim\mathrm{LKJ}(10)$ and multiplying $\mathbf{L}$ with the design matrix $\mathbf{X}$. For binary variables, we induce correlation with another variable using the following approach:
Data: $\mathbf{x}\in\mathbb{R}^{n}$, $r\in(-1,1)$
Result: $\mathbf{z}\in\{0,1\}^{n}$
$\mathbf{y}\sim\mathcal{N}_{n}(0,1)$;
$\mathbf{y}\leftarrow r\cdot\mathbf{x}+(1-r^{2})^{\frac{1}{2}}\cdot\mathbf{y}$;
$\mathbf{p}\leftarrow(1+e^{-\mathbf{y}})^{-1}$;
$\mathbf{z}\sim\mathrm{Bernoulli}(\mathbf{p})$;
@@ A.1 Predictors
A.2 Priors
@@
Data: $b\in\mathbb{N}$, $d\in\mathbb{N}$, $q\in\mathbb{N}$
Result: $\bm{\nu}_{\beta}\in\mathbb{R}^{b\times d}$, $\bm{\tau}_{\beta}\in\mathbb{R}^{b\times d}$, $\bm{\tau}_{\sigma}\in\mathbb{R}^{b\times q}$, $\bm{\tau}_{\varepsilon}\in\mathbb{R}^{b}$
$\bm{\nu}_{\beta}\sim\mathcal{U}_{b\times d}(-20,20)$;
$\bm{\tau}_{\beta_{0}}\sim\mathcal{U}_{b\times 1}(0.1,30)$;
$\bm{\tau}_{\beta}\sim\mathcal{U}_{b\times d-1}(0.1,20)$;
$\bm{\tau}_{\beta}\leftarrow[\bm{\tau}_{\beta_{0}},\bm{\tau}_{\beta}]$;
$\bm{\tau}_{\sigma}\sim\mathcal{U}_{b\times q}(0.1,10)$;
$\bm{\tau}_{\varepsilon}\sim\mathcal{U}_{b\times 1}(0.001,10)$;
@@
For the toy example, $\bm{\nu}_{\beta}$ is kept 0, $\bm{\tau}_{\beta}$ is bounded below $5$, and $\bm{\tau}_{\sigma}$ and $\bm{\tau}_{\varepsilon}$ are bounded below $1$.
@@ Appendix B Standardization
$\sigma_{0}^{*2}=\mathrm{Var}\left(\alpha_{i0}\right)+\mathrm{Var}\left(\sum_{k=1}^{q}\mu_{z_{k}}\alpha_{ik}\right)+2\cdot\mathrm{Cov}\left(\alpha_{i0},\sum_{k=1}^{q}\mu_{z_{k}}\alpha_{ik}\right),$
which is equivalent to summing up the covariance matrix of the random vector $\bm{\mu}_{z}\odot\bm{\alpha}_{i}$.
@@
Appendix C Data Representation and Embedding
Group-membership is represented implicitly by a separate tensor dimension, e.g. $\mathbf{X}$ has the shape $(batch,m,n,d)$. For PyTorch dataloader compatiblity, all tensors are zero-padded and corresponding masks are stored. To spread the learning signal evenly across the network, all slope-related variables are randomly permuted separately per regression dataset, using the same permutation for $\mathbf{X},\mathbf{Z},\bm{\beta},\bm{b}_{i}$, and $\mathbf{S}$.
@@ Appendix D Posterior Factorization
Let the joint distribution over all regression parameters and the data be
$p(\bm{\vartheta},\bm{\alpha},\mathbf{D}),$
where $\bm{\alpha}=\{\bm{\alpha}_{i}\}_{i=1,\ldots,m}$ and $\mathbf{D}=\{\mathbf{D}_{i}\}_{i=1,\ldots,m}$.
We can write the joint posterior as
$\frac{p(\bm{\vartheta},\bm{\alpha},\mathbf{D})}{p(\mathbf{D})}=p(\bm{\vartheta},\bm{\alpha}\,|\,\mathbf{D})=p(\bm{\vartheta}\,|\,\mathbf{D})\,p(\bm{\alpha}\,|\,\bm{\vartheta},\mathbf{D})=p(\bm{\vartheta}\,|\,\mathbf{D})\,\prod_{i=1}^{m}p(\bm{\alpha}_{i}\,|\,\bm{\vartheta},\mathbf{D}_{i}),$
@@
$\ell=\ell_{\Pi_{g},\Sigma_{g}}+\sum_{i=1}^{m}\ell^{(i)}_{\Pi_{l},\Sigma_{l}}\propto-\mathbb{E}_{\bm{\vartheta},\bm{\alpha},\mathbf{D}}\left[\log p_{\Pi_{g}}\left(\bm{\vartheta}\,|\,f_{\Sigma_{g}}(f_{\Sigma_{l}}(\mathbf{D}))\right)+\sum_{i=1}^{m}\log p_{\Pi_{l}}\left(\bm{\alpha}_{i}\,|\,\bm{\vartheta},f_{\Sigma_{l}}(\mathbf{D}_{i})\right)\right].$
@@
Appendix E Alternating Importance Sampling
@@
    $\log w_{i}\leftarrow\log p(\mathbf{D}|\bm{\vartheta}_{i})+\log p(\bm{\vartheta}_{i})-\log q(\bm{\vartheta}_{i}|\mathbf{D})$
@@
$p(\mathbf{D}|\bm{\vartheta})p(\bm{\vartheta})=\prod_{i=1}^{m}p\big(\mathbf{y}_{i}|\mathbf{X}_{i},\bm{\beta},\bm{\sigma}_{\alpha}^{2},\sigma_{\varepsilon}^{2}\big)p(\bm{\beta})p(\bm{\sigma}_{\alpha}^{2})p\big(\sigma_{\varepsilon}^{2}\big),$
@@
$p(\mathbf{D}|\bm{\vartheta})p(\bm{\vartheta})=\prod_{i=1}^{m}p\big(\mathbf{y}_{i}|\mathbf{X}_{i},\bm{\alpha}_{i},\bm{\beta},\sigma_{\varepsilon}^{2}\big)p(\bm{\alpha}_{i}|\bm{\sigma}_{\alpha}^{2})p(\bm{\sigma}_{\alpha}^{2})p(\bm{\beta})p\big(\sigma_{\varepsilon}^{2}\big).$
@@ Appendix F Result Figures
Descriptors are the same as in Figure 2 and Figure 3B.

### v2

@@ 4 Discussion
In this paper we present metabeta, a probabilistic transformer-based model that performs efficient approximate Bayesian inference for mixed-effects regression. We trained metabeta on simulated datasets with varying ranges for predictors, regression parameters, and outcomes. Most importantly, these datasets incorporate varying priors and we condition the model outputs on them, which not only amortizes the high computational costs encountered when using MCMC for parameter estimation, but also generalizes previous neural posterior estimation (NPE) techniques that are trained on a fixed prior. We show that our model has favorable and robust performance on in-distribution and out-of-distribution test sets, based on real hierarchical datasets practitioners care about. In each experiment, we compare the results of our model with Hamiltonian Monte Carlo (HMC) and variational inference (VI), the gold-standard methods for Bayesian mixed-effects regression, and show that metabeta generally approaches HMC in model fit, accuracy, and fidelity of credible intervals and outperforms VI in most. Most importantly, it does that at a small fraction of the time required for parameter estimation with conventional methods.
@@ 4.1 Limitations and Outlook
Our choice of model architecture trades of posterior expressivity for computation speed: Other normalizing flow methods like Neural Spline Flows (Durkan et al., 2019), Flow Matching (Wildberger et al., 2023), Conditional Diffusion (Chen et al., 2025; Reuter et al., 2025) or TarFlow (Zhai et al., 2025) offer more flexible posterior shapes, but posterior sampling is considerably more expensive than for Affine Coupling Flows, often involving numerical integration or solving a stochastic differential equation. The relative simplicity of affine coupling posteriors can be seen as implicit regularization, preventing overly irregular quantification of regression parameter uncertainty.
Each trained version of metabeta is currently tailored to the size of the regression problem in terms of the number of fixed effects ($d$) and the number of random effects ($q$). Together this collection of models acts like a single pretrained model, as each size can be pulled quickly from the repository for immediate deployment. That is, from the practitioners perspective it makes no difference if there is a single or multiple pretrained models for different regression problem sizes. Future work should strive towards a unified model.
Currently, the prior families are fixed. The parameters of the priors are concatenated to the summary vector $\mathbf{s}$ before being passed to the MLPs inside the normalizing flow. This approach could be generalized to varying prior families, whose identity can be embedded and simply added to the summary vector. We plan to eventually extend metabeta to even more flexible prior specification. Similarly, our framework is currently specialized on Bayesian linear mixed effects regression, but the required steps to generalized mixed-effects models are in parts small: Data simulation would require an additional response function around the linear term. The response function type could be passed to the model along with the priors. Extending importance sampling for non-linear cases is non-trivial, however. Finally, hierarchical NPE is well suited for mixed-effects regression with one grouping factor: Multiple parallel grouping factors would require non-trivial extensions to dataset summarization and integration of multiple summaries. However, it is conceptually straightforward to extend the framework to multiple nested grouping factors (e.g. states, schools, classrooms, students). Overall, these extensions are worthwhile avenues for future developments.
@@ 4.2 Conclusion
Our model brings Bayesian mixed-effects regression closer to practical usability in real-world applications. In its current form, it already enables rapid prototyping – practitioners can quickly test different model specifications and validate findings using conventional tools if needed. Our analyses highlight that metabeta is immediately applicable to such use cases. Looking ahead, we envision scaling our model to larger and non-linear regression problems. This would open the door to entirely new applications of Bayesian mixed-effects regression that are currently out of reach.
Acknowledgments
We thank Niki Kilbertus for his early-stage tips on model architecture, Fabian Scheipl for his advice on input standardization, Jakob Macke for his literature pointers, and Daniel Habermann for his valuable insights on data simulation and HMC sampling.
@@ Appendix A Dataset Simulation
All simulated or sampled datasets have $5\leq m\leq 30$ groups and each group has $10\leq n_{i}\leq 70$ observations, making $n$ range from $50$ to $2100$.
A.1 Synthetic Predictors
We sample $n_{i}$ observations of predictor $j$, $\mathbf{x_{ij}}$, from the following distributions: Normal, Student-$t$, continuous uniform, log-Normal, Bernoulli, negative binomial, and scaled Beta. All distributions have varying parameters and include random truncations. Correlation is induced by sampling $\mathbf{LL}^{\top}=\mathbf{R}\sim\mathrm{LKJ}(10)$ and multiplying $\mathbf{L}$ with the design matrix $\mathbf{X}$ (Lewandowski et al., 2009). For binary variables, we induce correlation with another variable using the following approach:
@@ A.1 Synthetic Predictors
[…]
1: Sample $\mathbf{y}\sim\mathcal{N}_{n}(\mathbf{0},\mathbf{I})$
[…]
4: Sample $\mathbf{z}\sim\mathrm{Bernoulli}(\mathbf{p})$
@@
A.2 Real Predictors
We use 271 real datasets from the PMLB (Romano et al., 2021) and SRM (Lichtenberg and Şimşek, 2017) benchmarks as additional sources for realistic predictors, and preserve hierarchical grouping structure when present. Existence of grouping structure is automatically checked by checking every non-continuous predictor for its number of unique values, as well as their spread. When such grouping factors are present, data is separately sampled per group, otherwise groups are randomly assigned. To further increase variability, we pass the sampled real data through Stochastic Gradient Langevin Dynamics (SGLD, Welling and Teh, 2011; Raginsky et al., 2017; Ma et al., 2024). This generates structurally equivalent data instead of just using subsets. The training sets receive a mix of synthetic and emulated predictors, the test sets receive only real data subsets. The in-distribution test sets rely on samples from the 271 datasets, the out-of-distribution sets rely on 7 additional hierarchical datasets not used in the training data.
A.3 Priors and Rescaling
@@
[…]
1: Sample $\boldsymbol{\nu}_{\beta}\sim\mathcal{U}_{b\times d}(-3,3)$
2: Sample $\boldsymbol{\tau}_{\beta}\sim\mathcal{U}_{b\times d}(0.01,3)$
3: Sample $\boldsymbol{\tau}_{\sigma}\sim\mathcal{U}_{b\times q}(0.01,3)$
4: Sample $\boldsymbol{\tau}_{\varepsilon}\sim\mathcal{U}_{b\times 1}(0.01,3)$
In a first forward pass, the standardized predictors and sampled parameters are projected to $\mathbf{y}$. Then all parameters (and their corresponding prior parameters) are rescaled such that $\mathbb{V}(\mathbf{y})=1$. This is without loss of generality, as posterior samples can trivially be brought back to the original scale of $\mathbf{y}$ by rescaling (see below). However, the advantage of this approach is that this covers a very wide range of potentially observable combinations of $\mathbf{X}$ and $\boldsymbol{\vartheta}$.
@@ Appendix B Standardization
$\sigma_{0}^{*2}=\mathbb{V}\left(\alpha_{i0}\right)+\mathbb{V}\left(\sum_{k=1}^{q}\mu_{z_{k}}\alpha_{ik}\right)+2\cdot\mathrm{Cov}\left(\alpha_{i0},\sum_{k=1}^{q}\mu_{z_{k}}\alpha_{ik}\right),$
which is equivalent to summing up the covariance matrix of the random vector $\boldsymbol{\mu}_{z}\odot\boldsymbol{\alpha}_{i}$.
@@
Appendix C Data Representation
Group-membership is represented implicitly by a separate tensor dimension, e.g. $\mathbf{X}$ has the shape $(batch,m,n,d)$. For PyTorch dataloader compatiblity, all tensors are zero-padded and corresponding masks are stored. To spread the learning signal evenly across the network, all slope-related variables are randomly permuted separately per regression dataset, using the same permutation for $\mathbf{X},\mathbf{Z},\boldsymbol{\beta},\boldsymbol{b}_{i}$, and $\mathbf{S}$.
@@ Appendix D Posterior Factorization
Let the joint distribution over all regression parameters and the data be $p(\boldsymbol{\vartheta},\boldsymbol{\alpha},\mathbf{D}),$ where $\boldsymbol{\alpha}=\{\boldsymbol{\alpha}_{i}\}_{i=1,\ldots,m}$ and $\mathbf{D}=\{\mathbf{D}_{i}\}_{i=1,\ldots,m}$. We can write the joint posterior as
$\frac{p(\boldsymbol{\vartheta},\boldsymbol{\alpha},\mathbf{D})}{p(\mathbf{D})}=p(\boldsymbol{\vartheta},\boldsymbol{\alpha}\,|\,\mathbf{D})=p(\boldsymbol{\vartheta}\,|\,\mathbf{D})\,p(\boldsymbol{\alpha}\,|\,\boldsymbol{\vartheta},\mathbf{D})=p(\boldsymbol{\vartheta}\,|\,\mathbf{D})\,\prod_{i=1}^{m}p(\boldsymbol{\alpha}_{i}\,|\,\boldsymbol{\vartheta},\mathbf{D}_{i}),$
@@
$\ell=\ell_{\Pi_{g},\Sigma_{g}}+\sum_{i=1}^{m}\ell^{(i)}_{\Pi_{l},\Sigma_{l}}\propto-\mathbb{E}_{\boldsymbol{\vartheta},\boldsymbol{\alpha},\mathbf{D}}\left[\log p_{\Pi_{g}}\left(\boldsymbol{\vartheta}\,|\,f_{\Sigma_{g}}(f_{\Sigma_{l}}(\mathbf{D}))\right)+\sum_{i=1}^{m}\log p_{\Pi_{l}}\left(\boldsymbol{\alpha}_{i}\,|\,\boldsymbol{\vartheta},f_{\Sigma_{l}}(\mathbf{D}_{i})\right)\right].$
@@
Appendix E Importance Sampling
@@
    $\log w_{i}\leftarrow\log p(\mathbf{D}|\boldsymbol{\vartheta}_{i})+\log p(\boldsymbol{\vartheta}_{i})-\log q(\boldsymbol{\vartheta}_{i}|\mathbf{D})$
@@
$p(\mathbf{D}|\boldsymbol{\vartheta})p(\boldsymbol{\vartheta})=\prod_{i=1}^{m}p\big(\mathbf{y}_{i}|\mathbf{X}_{i},\boldsymbol{\beta},\boldsymbol{\sigma}_{\alpha}^{2},\sigma_{\varepsilon}^{2}\big)p(\boldsymbol{\beta})p(\boldsymbol{\sigma}_{\alpha}^{2})p\big(\sigma_{\varepsilon}^{2}\big),$
@@
$p(\mathbf{D}|\boldsymbol{\vartheta})p(\boldsymbol{\vartheta})=\prod_{i=1}^{m}p\big(\mathbf{y}_{i}|\mathbf{X}_{i},\boldsymbol{\alpha}_{i},\boldsymbol{\beta},\sigma_{\varepsilon}^{2}\big)p(\boldsymbol{\alpha}_{i}|\boldsymbol{\sigma}_{\alpha}^{2})p(\boldsymbol{\sigma}_{\alpha}^{2})p(\boldsymbol{\beta})p\big(\sigma_{\varepsilon}^{2}\big).$
@@ Appendix F Result Figures
Descriptors are the same as in Figure 2. The first row corresponds to metabeta and the second to HMC. Additionally, posterior predictive distributions are plotted against the true outcomes $y_{ij}$ for two randomly chosen datasets from the test batch.